OpenStudy (anonymous):

Find: sqrt(21-12sqrt3)

OpenStudy (raden):

the formula to simplifying that is : sqrt( (a+b)-2sqrt(ab) ) = sqrt(a) - sqrt(b) with a > b first, setting 12 = 2 * 6 then it can be: